OpenStudy (anonymous):

Describe how to solve 5x^5-2x^3+x=0

OpenStudy (anonymous):

reply fast pl.

OpenStudy (anonymous):

\(\Large 5x^5-2x^3+x=0\\\Large x(5x^4-2x^2+1)=0\\\Large \Rightarrow x=0\ or\ 5x^4-2x^2+1=0\) Let \(\Large t=x^2\) \(\Large \Rightarrow 5t^2-2t+1=0 \) this equation has no solution So, just have 1 solution:\(\Large x=0\)

OpenStudy (anonymous):

thatnk one more

OpenStudy (anonymous):

solve 8sqrt(x)+240